OpenStudy (anonymous):

You have 80 yards of fencing to enclose a rectangular region.find the dimensions of the rectangle that maximize the enclose area.what is the maximize area?

OpenStudy (anonymous):

I need all the steps!

OpenStudy (anonymous):

a+a+b+b=80 2a+2b=80 a+b=40 a=40-b area=A=a*b A=(40-b)*b A=40b-b^2 To get a maximum value, take the derivative of A with respect to b, and set it to zero. dA/db=40-2b 0=40-2b 2b=40 b=20 a=40-b a=40-20 a=20 20 yards by 20 yards

OpenStudy (anonymous):